Urgent please help!!!
-1+41log3(-3x)=7
Solve please!!!!

Answer by Edwin McCravy(6935)

Urgent please help!!!

 -1 + 41log3(-3x) = 7

Isolate the log term by adding 1
to both sides:

     41log3(-3x) = 8

Divide both sides by 41

       log3(-3x) = 8/41

Use this rule of logarithms:

The logarithm equation logBA = C can be 
rewritten equivalently as the exponential
equation A = BC

           -3x = 38/41

Divide both sides by -3

             x = (38/41) ÷ (-3) 

             x = -(38/41)÷(31)

Subtract exponents of 3:

             x = -38/41-1

             x = -38/41-41/41

             x = -3-33/41

             x = -1/(333/41) 
  
             x = -.413024273, approximately

Edwin